# MemANNS: Enhancing Billion-Scale ANNS Efficiency with Practical PIM Hardware

### v1

@@ 3. System Overview
In viewing the limitations of existing ANNS solutions, this paper presents MemANNS, an efficient system that accelerates the ANN search using UPMEM PIM hardware. Figure 5 shows an overview of our system. The offline phase is executed on the host CPU to prepare and transfer data to DPUs for online processing. The cluster filtering stage of the online phase is executed on the host CPU to filter clusters and the rest stages are executed on the DPUs to utilize their high aggregated memory bandwidth. Optimizations have been proposed to improve the efficiency and performance of the offline and online phases of the IVFPQ algorithm separately.
Offline Optimizations. Recall that the offline phase uses IVF and PQ techniques to encode data points into smaller sizes. The encoded data are subsequently transferred to PIM for further processing. We propose two techniques for the offline phase. First, we propose a PIM-aware data placement technique which strategically distributes encoded vectors across hundreds of DPUs, ensuring a balanced workload for optimized performance (tackling Challenge 1). Second, we design a co-occurrence aware vector encoding strategy that takes item co-occurrences into account to reduce the memory accesses during online search, thereby alleviating the workload pressure on PIM (tackling Challenge 3).
Online Optimizations. The online phase of IVFPQ takes a batch of queries as input, and conducts four stages to search for the nearest neighbors in the encoded vectors for the query vectors. We accelerate the online search process by deploying the most memory-intensive stages onto PIM to utilize its high aggregated memory bandwidth. We mainly propose the following optimizations to fully utilize PIM resources. First, we propose a balanced query scheduling approach to map the search workloads onto different DPUs, considering the placement of encoded vectors decided at offline phase. This scheduling approach works together with the data placement to achieve workload balance. Second, we propose two PIM resource management techniques, including thread scheduling and memory management, to fully explore the capabilities of PIM (tackling Challenge 2). Third, to prevent the top-k identification stage becoming a new bottleneck in PIM, as illurstrated in Figure 1 for GPUs, we propose a top-k pruning technique to improve its performance.
@@ 4. Technical Details
This section presents the technical details of the optimization techniques in MemANNS, and explains how they tackle the challenges mentioned in Section 2.3.
@@ 4.1. PIM-Aware Workload Distribution
The workload distribution task includes two sub-tasks, namely the data placement at offline time and query scheduling at online phase. The data placement task tries to distribute the encoded data points across DPUs as balanced as possible, considering the variant cluster access frequencies and cluster sizes; while the query scheduling task tries to map the filtered clusters of all queries in the batch to different DPUs to actually balance the workload at runtime.
Offline Data Placement. The goal of the data placement is two-fold, namely avoiding inter-DPU/CPU-DPU data transfers and balancing the online data searching workload (i.e., the amount of memory accesses to the MRAM).
Considering that the online phase processes data according to clusters, a natural idea is to place an entire cluster in the same DPU, and maintain a local top-k list on each DPU for a query. The local top-k are then transferred to the CPU to generate the final top-k. By placing an entire cluster in a single DPU, we reduce the amount of partial top-k transmitted for a query, alleviating the communication bottleneck between DPUs and the host CPU. Following this idea, we make data placement decisions based on clusters. Algorithm 1 shows the flow of our data placement approach.
The workload of a cluster is mainly contributed by the memory-intensive distance calculation stage, which performs excessive memory search in LUTs to calculate distances between a query $q$ and all search points in the cluster. Thus, we can estimate the workload of a cluster $i$ as follows. Let $s_{i}$ denote the size of cluster $i$ and $f_{i}$ denote the frequency with which cluster $i$ is used to process queries. The workload of cluster $i$ can be represented as $w_{i}=s_{i}*f_{i}$. Given $n$ DPUs and a batch of queries, the balanced memory accesses per DPU that we need to achieve is $\overline{W}=\frac{1}{n}\sum_{i}s_{i}\times f_{i}$, where $f_{i}$ is derived from a predictor based on historical query data. The basic idea of achieving workload balance is to distribute data clusters across DPUs so that the workload on each DPU approximates $W$ as closely as possible.
@@
1:$ndpu$: the number of available DPUs; $s_{i}$: #vectors in cluster $i$;
@@
$MAX\_DPU\_SIZE$: the maximum number of vectors each DPU can have;
2:$dpu\_id$;
3:$d\_id\leftarrow ndpu$, $thld\leftarrow 1.0$ and $rate\leftarrow 0.02$;
4:$ncpy=\lceil{s_{i}*f_{i}}/{\overline{W}}\rceil$; $\triangleright$ $ncpy$ denotes #DPUs that cluster $i$ is distributed onto;
5:$w_{i}={s_{i}*f_{i}}/{ncpy}$; $\triangleright$ $w_{i}$ denotes the per DPU workload of cluster $i$;
@@
8:       $dpu\_id.\text{push}(d\_id)$, $ncpy\leftarrow ncpy-1$;
9:    $d\_id\leftarrow(d\_id+1)\mod ndpu$;
10:    if $d\_id==ndpu$ then $\triangleright$ no suitable DPU found;
11:       $thld\leftarrow thld+rate$; $\triangleright$ enlarge threshold to loosen the workload balance constraint;
Due to the skewness in cluster popularity and size, some clusters may have workloads significantly exceeding $W$. Thus, we create $ncpy$ copies of these high-demand clusters (Line 2) and distribute them across different DPUs. When such clusters are required, we schedule the queries to different copies to balance the workload. When placing each copy (ordered according to workload from high to low), we iterate over all DPUs with sufficient capacity and select the one that does not exceed $W$ after the cluster is placed (Line 5). If no suitable DPU is found, we relax the workload balance constraint by adjusting the parameter $thld$ (Line 9) and restart the search. At the end of this algorithm, each cluster is placed on a specific DPU (or a list of $ncpy$ DPUs).
Algorithm 1 optimizes the placement for one cluster, while it is also important to consider the cluster co-location problem according to their distances. We observe from real-world datasets that, clusters selected for a query are typically situated in close proximity. For example, as shown in Figure 7, the three centroids in the red circle are likely to be selected concurrently for the query. Note that, each cluster will generate a partial top-k, which will be transmitted to the CPU to generate the final top-k. As CPU-DPU communications go through the memory bus, which is much slower than intra-DPU data exchange, we need to reduce the amount of partial top-k generated during query processing. Thus, when a DPU is selected for cluster $i$, we will continue to assign nearby clusters to the same DPU, based on inter-cluster distances, until the workload of the DPU reaches $W$. This will allow the partial top-k results of co-located clusters to be aggregated locally to dramatically reduce communication overhead.
Figure 7 shows the workload distribution and memory usage of the SIFT1B dataset optimized using our data placement approach. For detailed setting, please refer to Section 5. The results show that, with the above optimizations, we are able to achieve a balanced workload across DPUs and the memory size used on each DPU does not vary significantly.
Figure 6. Inter-cluster Distance
Figure 7. Workload distribution across DPUs
Online Query Scheduling. At online phase, the cluster filtering stage selects $nprobe$ clusters for each query. This stage only compares the centroid vectors with the query vectors for filtering, thus is relatively light-weighted compared to the other three stages. We execute this stage and the query scheduling algorithm on the host CPU and send the scheduling results to PIM for actual searching.
Since some popular clusters are replicated and distributed across DPUs, the query scheduling algorithm selects the appropriate replica of the cluster for a query to ensure workload balance. This can be considered as a load balancing bin-packing problem with constraints on which bins each item can fit into. We design a greedy algorithm to solve the problem as shown in Algorithm 2. Specifically, we first schedule the clusters with only one replica (Line 4-5). Next, we update the workload on each DPU (Line 6) and continue to schedule the clusters with more than one replicas (Line 8-14). We order these clusters in descending order according to cluster sizes, and iteratively select the DPU with the least workload for each cluster. The algorithm returns a vector named $Assigned$, where the $i$-th element contains the list of $\langle q\_id,c\_id\rangle$ pairs scheduled to execute on DPU $i$. According to $Assigned$, we send the corresponding $q-c$ vectors to the DPUs for execution. Algorithm 2 runs at online phase, with a complexity of $O(|Q|\times nprobe)$. Both $|Q|$ and $nprobe$ are very small compared to the billion-scale data sizes, making the overhead of this algorithm negligible.
@@
2:$Assigned$: the query scheduling result;
3:$W[i]\leftarrow 0$ for $i\in\{0,\ldots,ndpu-1\}$; $\triangleright$ $W[i]$ denotes the workload on DPU $i$
@@
4.2. PIM Resource Management
In the last subsection, the PIM-aware workload distribution has promised balanced workloads across DPUs. In this subsection, we discuss how to efficiently manage the limited resources within each DPU to improve system performance. Specifically, we optimize the thread scheduling and memory management on PIM to better meet application needs.
Thread Scheduling. The UPMEM architecture provides a parallel structure between DPUs, and each DPU is equipped with up to 24 hardware threads and a 14-stage pipeline, effectively mitigating the impact of memory access latency. To efficiently utilize the multi-threads, we can explore hardware parallelism at various levels: between queries, among different clusters of a single query, or within a cluster.
Note that multiple threads operating on the same DPU share the limited resources of 64KB WRAM. Since accessing data from WRAM is much faster than from MRAM, we try to fit all data needed during thread execution in WRAM, including the codebook, LUT and encoded data in selected clusters. The codebook size can be estimated as $D\times 256$, where $D$ represents the dimension of data points, amounting to 32KB for the SIFT dataset. The LUT size can be estimated as $M\times 256\times sizeof(uint16)$, which is 8KB when the dimension of encoded points $M$ is 16. If we pursue parallelism at the query or cluster level, the WRAM space will be over-utilized since the combined size of the LUTs for more than four clusters will easily exceed 64KB. Given this hardware constraint, we implement multi-threading within each cluster.
@@ 4.2. PIM Resource Management
Each DPU is assigned with multiple clusters according to the query scheduling results of Algorithm 2. On each DPU, clusters are processed sequentially while multiple threads execute concurrently to accelerate the processing of a single cluster. Figure 9 illustrates our multi-threading strategy. Specifically, multiple threads split the LUT construction workload by concurrently reading the codebook $B$ from MRAM to WRAM to create a segment of the LUT. After the LUT is constructed, the partial sum of high-frequency co-occurrence combinations are computed using the LUT (details introduced in Section 4.3). The partial sums and the LUT will be combined to construct the full LUT. Once the full LUT is constructed, multiple threads concurrently read the encoded points $E$ from MRAM to WRAM to calculate the partial distances between the query and search points. Pipelining is enabled to overlap MRAM reads and distance calculations to hide the high latency. Each thread maintains a thread-local priority queue (PQ), implemented as a max heap with a size of $k$, to store the local top-k partial distances. After a thread finishes the partial top-k insertion, it moves on to process the next cluster.
@@
Memory Management. Due to the significant bandwidth difference between WRAM and MRAM, we aim to maximize the data stored in WRAM. The UPMEM DPU lacks a Memory Management Unit (MMU) to virtualize its physical memory, and the DPU is constrained to the limited physical memory capacity. Therefore, we propose a WRAM reuse strategy specifically tailored to the query processing stages on PIM. This strategy enables multiple threads to operate in a larger address space than the physical WRAM space available.
Figure 9 illustrates the reuse strategy using the SIFT dataset as an example. During the LUT construction stage, we use multiple threads to read the codebooks (32KB) and compute the corresponding entries in the LUT (8KB). Once the complete LUT is obtained, we use the same capacity to construct the sum of the combinations (8KB). Since the codebooks will not be used in the following phases, they can be safely overwritten to conserve space. As a result, the total space required for the codebooks and the full LUT is only 48KB. Once the full LUT is obtained, we proceed to the distance calculation stage, which is the most time-consuming stage due to the costly encoded point fetching from MRAM to WRAM. To optimize this process, we reuse the WRAM space occupied by codebooks to allow more threads to load encoded data points concurrently. In the example shown in Figure 9, we utilize 16 threads consuming 32KB of memory.
The read latency of MRAM does not increase linearly with size, as shown in Figure 9. The MRAM read latency increases slowly as the data size grows from 8B to 256B. Beyond 256B, the latency rises nearly linearly. This suggests that smaller MRAM read sizes (under 256B) yield greater benefits. Larger MRAM reads can consume significant WRAM space with minimal returns. We explore an optimal MRAM read size (buffer size) to enhance overall efficiency.
@@ 4.3. Co-occurrence Aware Encoding
The limited computation and storage resources on a single DPU motivate us to prune the workload needed on DPUs to further improve ANNS efficiency. Existing work have proposed various workload pruning techniques using distance bound (André et al., 2016; Gao and Long, 2023) and machine learning (Tian et al., 2024). These methods can be useful to MemANNS, but are still not good enough.
We observe that items in the encoded points are indices pointing to the codebooks, thus resulting in limited values that typically range from 0 to 255. This allows for certain item combinations to co-occur. Figure 10 shows the maximum co-occurrence frequency of item combinations with lengths 3, 4 and 5 in the SIFT1B dataset. For example, when the length is 3, there could be $256^{3}$ possible item combinations. We count the occurrences of each combination in the encoded points and Figure 10 shows the maximum counts at different lengths. Notably, shorter combinations tend to have a higher frequency of co-occurrence. For example, the most frequently co-occurring combination of length 3 was found in 5.7% of all encoded points. By caching the partial sums of these high‐frequency combinations at offline, we can reduce the memory accesses and computations required during the distance calculation stage at runtime, hence greatly alleviating the performance bottleneck.
Figure 10. Max. freq. of comb. with different lengths
Avg. Length
Reduction
Time
Reduction
0 0
0.25 18%
0.5 36%
0.75 54%
Table 1. Higher code length reduction leads to higher query search time reduction
Utilizing the above observation, we propose a co-occurrence aware data encoding. Specifically, we adopt a graph-based approach (Ye et al., 2023) which builds an Item Co-occurrence Graph (ICG) to track item co-occurrences and apply a clustering algorithm to identify frequently accessed combinations of varying lengths. For each cluster, we use this algorithm (Ye et al., 2023) to select $m$ most frequent combinations of length 3. Combinations with longer lengths can be selected if given larger cache size (WRAM size). The parameter $m$ is also decided by WRAM size and by default, we set it to 256. Our goal is to cache the partial sum of these combinations to avoid repeated computations during online search. However, the distances are unknown at offline phase and the partial sum cannot be pre-computed. Instead, we propose to reserve a buffer space in WRAM to store the partial sum obtained after the LUT construction stage. We pre-arrange the buffer space and generate the memory address for each partial sum. The cache address will be used to re-encode the IVFPQ encoded points. Figure 11 gives a detailed example.
First, as shown on the top of the figure, we use ICG to identify three co-occurrence item sets, namely (1, 15, 26), (79, 25, 77) and (2, 14, 31). Note that, different from existing work (Ye et al., 2023) which only considered the occurrence of certain items, we must also take into account the positions where these occurrences happen. For example, when utilizing the cached partial sum of the combination (1, 15, 26), it is crucial that these items appear in the columns (0, 1, 2) respectively to make the sum meaningful. Second, we calculate the cache address for the partial sum of all possible combinations in the co-occurrence sets. Third, we use the generated cache address to re-encode an encoded vector, which has 16 dimensions and each dimension is a $uint8$ value in the 0-255 range. This vector contains the high-frequency combinations (1, 15, 26), (79, 25, 77) and (14, 31) on the correct positions. According to the pre-computed cache address, we can find the partial sum of the three combinations at 0x00111, 0x01111 and 0x10110 respectively. Thus, in the new encoded vector, we use the address of the partial sum to replace the original items in the vector. Specifically, the new encoded vector contains three parts. The first digit represents the total length of the vector, followed by the original code which are not identified as high-frequency co-occurred items. These codes are stored as $uint16$ to include the column index information. The combination codes are appended in the end, represented using $uint16$ cache addresses of partial sums. The entire process described above is executed during the offline phase.
@@
During online phase, we calculate the partial sums using constructed LUT and store the sums in WRAM according to pre-arranged layout. For example, for combination (1, 15, 26) on position (0, 1, 2), the partial sum is calculated as $LUT[1+0\times 256]+LUT[15+1\times 256]+LUT[26+2\times 256]$, and stored at address 0x00111. By reusing the cached partial sum, we can greatly reduce the computation and memory accesses needed during distance calculation stage.
@@
Note that, during partial sum calculation, we need multiplication operations to access the LUT (e.g., $15+1\times 256$). However, existing work (Gómez-Luna et al., 2022) shows that multiplication operations on PIM are less efficient. To mitigate this issue, as shown in Figure 11, we further modify the new encoded vector by converting the original code and combination code into direct addresses. The direct address of the original code can be calculated using the code and its column index. We store partial sums in a contiguous memory space after LUT. Thus, the direct address of a combination code can be calculated using the total size of LUT and the code. For example, for combination code 7, it is converted to $16\times 256+7=4103$.
@@
In the example, our co-occurrence aware encoding reduces the length of the encoded vector from 16 to 12, achieving a 25% reduction in length. Our empirical studies indicate that, a higher length reduction rate corresponds to a greater decrease in distance calculation time, as demonstrated in Table 1. In our evaluations, we implement this new encoding strategy only when the average length reduction exceeds 50% to avoid negatively affecting the compression rate.

### v2

@@ 3. System Overview
This paper presents UpANNS, an efficient system that accelerates ANN search using UPMEM PIM hardware. UpANNS incorporates four optimizations, targeting specific challenges mentioned above.
Opt1: PIM-Aware Workload Distribution. To mitigate skewed cluster access patterns (challenge 1), we dynamically replicate and distribute IVF clusters across DPUs during offline preprocessing. High-demand clusters are replicated to avoid hotspots, while spatially proximate clusters are co-located on the same DPU to minimize costly inter-DPU communication. This ensures balanced memory accesses and maximizes the utilization of UPMEM’s 7.2 TB/s aggregated bandwidth.
Opt2: PIM Resource Management. With UPMEM’s unique hardware architecture (challenge 2), we propose two optimizations to fully utilize its limited resources: 1) Thread scheduling: we utilize 24 hardware threads per DPU to overlap MRAM transfers with computations via a 14-stage pipeline; 2) Memory management: we carefully reuse WRAM buffers across IVFPQ stages (e.g., codebook, LUT to encoded vectors) to minimize MRAM accesses (Figure 7).
Opt3: Co-occurrence Aware Encoding. Targeting challenge 3, we identify frequent element combinations in encoded vectors and precompute partial sums during offline processing. This can greatly reduce online memory accesses without compromising accuracy.
Opt4: Top-K Pruning. We further address challenge 3 by eliminating redundant computations. For the bottlenecked top-k selection stage, we combine thread-local heaps with early termination, which enables skipping 68% of redundant comparisons and accelerates the final stage by 3.1x.
Figure 5 shows the optimized workflow of IVFPQ. During offline phase, clusters are encoded (Opt3), replicated and distributed (Opt1). During online phase, queries are scheduled to balanced DPUs (Opt1), and processed using WRAM-optimized pipelines (Opt2) and pruned top-k selection (Opt4).
@@ 4.1. PIM-Aware Workload Distribution
To address the severe workload imbalance caused by skewed cluster access patterns, UpANNS integrates two synergistic strategies: offline data placement and online query scheduling. The data placement task selectively replicates and distributes clusters across DPUs based on their access frequency patterns and sizes, enhancing system flexibility for diverse query workloads. Meanwhile, the query scheduling task efficiently maps filtered clusters from batch queries to appropriate DPUs, dynamically balancing workloads at runtime.
4.1.1. Data Placement.
The offline data placement strategy optimizes cluster distribution across DPUs based on three three key insights: (1) cluster-level locality preservation reduces partial result transfers, (2) workload skew necessitates adaptive replication, and (3) spatial proximity between frequently co-accessed clusters enables communication optimizations. Our data placement strategy carefully considered these three insights to achieve optimal performance and the key steps are shown in Algorithm 1.
Firstly, entire clusters are placed on a single DPU to avoid partial result transfers on the bottlenecked communication between DPUs and the host CPU. For a cluster $i$, its workload is mainly contributed by the memory-intensive distance calculation stage, which performs excessive memory search in LUTs to calculate distances between a query $q$ and all search points in the cluster. Thus, we can estimate the workload as $W_{i}=s_{i}*f_{i}$, where $s_{i}$ is the cluster size and $f_{i}$ is its historical access frequency. Given $n$ DPUs and a batch of queries, the balanced memory accesses per DPU that we need to achieve is $\overline{W}=\frac{1}{n}\sum_{i=1}^{m}W_{i}$, where $m$ is the number of clusters.
@@
1:ndpundpu: the number of available DPUs; $s_{i}$: #vectors in cluster $i$;
@@
MAX_DPU_SIZEMAX\_DPU\_SIZE: the maximum number of vectors each DPU can have;
2:dpu_iddpu\_id;
[…]
@@
8:   $count\leftarrow 0$;
9:   $dpu\_id.\text{push}(d\_id)$, $ncpy\leftarrow ncpy-1$;
10:  else
11:   $count\leftarrow count+1$;
12:  $d\_id\leftarrow(d\_id+1)\mod ndpu$;
13:  if $count==ndpu$ then $\triangleright$ no suitable DPU found;
14:   $thld\leftarrow thld+rate$; $\triangleright$ enlarge threshold to loosen the workload balance constraint;
Secondly, cluster access frequency is highly skew, as shown in Figure 4. Thus, some clusters may have workloads significantly exceeding $\overline{W}$. We create ncpyncpy copies for these high-demand clusters (Line 2) and distribute them across multiple DPUs. During placement, replicas are assigned to DPUs with the least residual capacity, iteratively increasing the acceptable workload threshold thldthld until all replicas are placed (Lines 5–12). At the end of this algorithm, each cluster is placed on a list of ncpyncpy DPUs.
Thirdly, clusters selected for a single query are typically in close proximity (e.g., neighboring centroids). We co-locate such clusters on the same DPU, which allows local aggregation of partial top-k results during multi-cluster queries, dramatically cutting inter-DPU communication overhead.
@@
2:AssignedAssigned: the query scheduling result;
[…]
@@
4.1.2. Query Scheduling.
At runtime, the host CPU assigns queries to DPUs using a greedy scheduling algorithm (Algorithm 2) to achieve workload balance dynamically. Each query selects nprobenprobe clusters by comparing the centroid vectors with the query vector, thus is relatively light-weighted compared to the other online stages. We execute cluster filtering and query scheduling on the host CPU and send the scheduling results to PIM for query searching.
Since some popular clusters are replicated and distributed across DPUs, the query scheduling algorithm dynamically selects the appropriate cluster replicas to ensure workload balance. Specifically, we first schedule the clusters with only one replica (Line 4-5). Next, we update the workload on each DPU (Line 6). Clusters with multiple replicas are ordered by size (descending) and assigned to the least-loaded DPU (Line 8-14). This ensures larger clusters do not overload individual DPUs. Algorithm 2 runs at runtime with a complexity of $O(|Q|\times nprobe)$, where $Q$ is the set of queries. Since both $|Q|$ and nprobenprobe are significantly smaller than billion-scale datasets, the overhead of this algorithm is negligible.
Due to hardware limitations preventing direct DPU-DPU communication, PIM faces challenges in handling changing query patterns. UpANNS targets applications such as retrieval-augmented LMs and recommendation systems, where query patterns typically change regularly (e.g., every few days (Jiang et al., 2023a)) and incrementally. To handle such changes, UpANNS implements an adaptive approach: (1) adjusting the number of cluster copies (Section 4.1.1) to accommodate minor query pattern changes in the short term, and (2) performing full data relocation to address major pattern shifts over longer time periods. Furthermore, our experiments demonstrate that the PIM hardware achieves significantly higher query throughput than CPU implementations even with random data distribution, highlighting the substantial advantages of PIM architectures.
4.2. PIM Resource Management
To maximize UPMEM’s limited per-DPU resources (small WRAM cache and 350 MHz DPU cores), UpANNS employs two synergistic optimizations: thread scheduling and memory management. These ensure full utilization of PIM’s parallel compute and memory bandwidth while mitigating latency bottlenecks.
4.2.1. Thread Scheduling.
The UPMEM architecture provides a parallel structure between DPUs. Further, each DPU’s 24 hardware threads and 14-stage pipeline enable fine-grained parallelism, but WRAM constraints (64KB) restrict concurrent cluster processing.
To efficiently utilize the multi-threads, we can explore hardware parallelism at various levels: inter-queries, between different clusters of a single query, or intra-cluster. Note that multiple threads operating on the same DPU share the limited 64KB WRAM. Since accessing data from WRAM is much faster than from MRAM, we try to fit all data needed during thread execution in WRAM, including the codebook, LUT and encoded data in selected clusters. The codebook size can be estimated as $D\times 256$, where $D$ represents the dimension of data points, amounting to 32KB for the SIFT dataset. The LUT size can be estimated as $M\times 256\times sizeof(uint16)$, which is 8KB when the dimension of encoded points $M$ is 16. If we pursue parallelism at the query or cluster level, the WRAM space will be over-utilized since the combined size of the LUTs for more than four clusters will easily exceed 64KB. Given this hardware constraint, we implement intra-cluster parallelism, where threads collaborate on a single cluster at a time.
On each DPU, multiple threads execute concurrently to accelerate the processing of a single cluster. Figure 7 illustrates our multi-threading strategy. Specifically, in the LUT construction stage, threads concurrently fetch codebook segments from MRAM into WRAM, constructing partial lookup tables (LUTs). For $M=16$ subvectors, this reduces LUT build time by 4.8x compared to single-threaded execution. After the LUT is constructed, the partial sum of high-frequency co-occurrence combinations are computed using the LUT (details introduced in Section 4.3). The partial sums and the LUT will be combined to construct the full LUT. Once the full LUT is constructed, multiple threads concurrently read the encoded points $E$ from MRAM to WRAM to calculate the partial distances between the query and search points. Pipelining is enabled to overlap MRAM reads and distance calculations to hide the high latency. Each thread maintains a thread-local priority queue (PQ), implemented as a max heap with a size of $k$, to store the local top-k partial distances. After a thread finishes the partial top-k insertion, it moves on to process the next cluster.
@@ 4.2. PIM Resource Management
4.2.2. Memory Management.
Due to the significant bandwidth difference between WRAM and MRAM, we aim to maximize the data stored in WRAM. The UPMEM DPU lacks a Memory Management Unit (MMU) to virtualize its physical memory, and the DPU is constrained to the limited physical memory capacity. Therefore, we propose a WRAM reuse strategy specifically tailored to the query processing stages on PIM. This strategy enables threads to operate in a larger address space than the physical WRAM space available.
Figure 7 illustrates the reuse strategy using the SIFT dataset as an example. During the LUT construction stage, we use multiple threads to read the codebooks (32KB) and compute the corresponding entries in the LUT (8KB). Once the complete LUT is obtained, we use the same capacity to construct the sum of the combinations (8KB). Since the codebooks are not used in the following stages, they can be safely overwritten to conserve space. As a result, the total space required for the codebooks and the full LUT is only 48KB. Once the full LUT is obtained, we proceed to the costly distance calculation stage, which needs to fetch encoded point from MRAM to WRAM. To optimize this process, we reuse the WRAM space occupied by codebooks to allow more threads to load encoded data points concurrently. In the example shown in Figure 7, we utilize 16 threads consuming 32KB of memory.
The read latency of MRAM does not increase linearly with size, as shown in Figure 7. The latency increases slowly as data size grows from 8B to 256B and increases almost linearly beyond 256B. This suggests that smaller MRAM read sizes (under 256B) yield greater benefits. Larger MRAM reads can consume significant WRAM space with minimal returns. We explore an optimal MRAM read size (buffer size) to enhance overall efficiency.
@@ 4.3. Co-occurrence Aware Encoding
The limited computational and memory resources of individual DPUs necessitate aggressive workload reduction without sacrificing accuracy. Existing works have proposed various workload pruning techniques using distance bounds (André et al., 2016; Gao and Long, 2023) and machine learning (Tian et al., 2024). These methods are complementary to UpANNS.
Encoded points are indices pointing to codebooks, resulting in a limited value range (typically within [0, 255]). This property allows frequent co-occurring element combinations. For instance, in the SIFT1B dataset, the triplet (1, 15, 26) appears in 5.7% of vectors. During distance calculation stage (Figure 2(c)), each element in an encoded vector triggers a LUT lookup, summing to compute the final distance. By caching the partial sums of the high-frequency combinations, we can reduce redundant memory accesses and computations during distance calculation at runtime, hence improving the querying performance.
@@
Leveraging the above observations, we propose a co-occurrence aware data encoding strategy. Specifically, we first identify high‐frequency co‐occurring element combinations in encoded points using an Element Co‐occurrence Graph (ECG). The ECG is constructed such that nodes represent elements and edges represent co‐occurrence relationships, with edge weights corresponding to co‐occurrence frequencies. For each cluster, we select the top-$m$ most frequent combinations of length 3. Longer combinations can be selected if a larger cache size (e.g., WRAM capacity) is available. The parameter $m$ is determined by the WRAM size; by default, we set mm=256. Our goal is to cache the partial sums of these combinations to avoid redundant computations during online searches. However, since distances are unknown during the offline phase, partial sums cannot be precomputed. To address this, we reserve buffer space in WRAM to store partial sums generated after the LUT construction stage. We preallocate this buffer space and assign memory addresses to each partial sum. These addresses are then used to re‐encode the IVFPQ‐encoded points. Figure 8 illustrates this process with a detailed example.
First, as shown on the top of the figure, we identify three co-occurrence element sets, namely (1, 15, 26), (79, 25, 77) and (2, 14, 31). Note that, we must also take into account the positions where these occurrences happen. For example, when utilizing the cached partial sum of the combination (1, 15, 26), it is crucial that these elements appear in the columns (0, 1, 2) respectively to make the sum meaningful. Second, we calculate the cache address for the partial sum of all possible combinations in the co-occurrence sets. Third, we use the generated cache address to re-encode an encoded vector, which has 16 dimensions and each dimension is a $uint8$ value in the 0-255 range. This vector contains the high-frequency combinations (1, 15, 26), (79, 25, 77) and (14, 31) on the correct positions. According to the pre-computed cache address, we can find the partial sum of the three combinations at 0x00111 (7), 0x01111 (15) and 0x10110 (22) respectively. Thus, in the new encoded vector, we use the address of the partial sum to replace the original elements in the vector. Specifically, the new encoded vector contains three parts. The first digit represents the total length of the vector, followed by the original codes which are not identified as high-frequency co-occurred elements and the combination codes which record the cache address of the partial sum.
@@
During online phase, we calculate the partial sums using constructed LUT and store the sums in WRAM according to pre-arranged layout. For example, for combination (1, 15, 26) on position (0, 1, 2), the partial sum is calculated as $LUT[1+0\times 256]+LUT[15+1\times 256]+LUT[26+2\times 256]$, and stored at address 0x00111. By reusing the cached partial sum, we can greatly reduce the computation and memory accesses during distance calculation stage.
@@
Note that, during partial sum calculation, we need multiplication operations to access the LUT (e.g., $15+1\times 256$). However, existing work (Gómez-Luna et al., 2022) shows that multiplication operations on PIM are less efficient. To mitigate this issue, as shown in Figure 8, we further modify the new encoded vector by converting the original code and combination code into direct addresses. For the original code, we calculate the direct address of the LUT by $256\times idx+original\_code$. For the combination code, we directly use the cache address with an offset of the LUT size ($256\times 16$). The length of the new encoded vector is stored in the second digit, which is 11 in this example. The reason for storing in the second digit is to ensure that the length of the new encoded vector is no more than the original length. If the new vector doesn’t have a high frequency combination, the length doesn’t need to be stored and the second digit is larger than $256\times 1$. Otherwise, the length is stored in the second digit and the second digit is less than $256\times 1$. We could identify the second digit to determine the length of the encoded vector.
@@
In the example, our co-occurrence aware encoding reduces the length of the encoded vector from 16 to 12, achieving a 25% reduction in length. Our empirical studies indicate that, a higher length reduction rate corresponds to a greater decrease in distance calculation time, as demonstrated in Figure 14.